# MSGNav: Unleashing the Power of Multi-modal 3D Scene Graph for Zero-Shot Embodied Navigation

### v3

@@ 1 Introduction
Embodied navigation is a fundamental capability of robotic agents, enabling them to accomplish diverse tasks in realistic environments [45, 52]. Such tasks are typically goal-oriented, with targets defined by objects, natural language descriptions, or images [43]. Recently, the challenging open-vocabulary embodied navigation task [55] has gained increasing attention, as real-world agents are required to generalize to a wide and unconstrained range of object categories [45].
Previous RL-based embodied navigation methods [28, 33] suffer from poor generalization and a large sim-to-real gap [42]. Inspired by rapid advances in pre-trained large model techniques [14, 1], zero-shot embodied navigation approaches have emerged at this historic moment. Zero-shot navigation system does not require any training or fine-tuning before being applied to real-world scenarios. These approaches [41, 15, 47, 56] substantially reduce training time and computational cost while enabling open‑vocabulary navigation, thereby facilitating deployment in real‑world applications.
Notably, a series of methods [36, 38, 30] that build explicit 3D scene graphs from observations and leverage LLMs to drive exploration have shown promising performance on standard goal-oriented navigation benchmarks [31, 2]. However, traditional 3D scene graphs overly abstract 3D object relations into simple textual labels (e.g., “top”, “beside”), which seriously impedes progress on zero-shot embodied navigation. As illustrated in Fig. 1, the extreme abstracted 3D object-relation in traditional 3D scene graphs introduces three main limitations:
@@
Figure 2: Performance comparisons between our MSGNav and other existing methods for embodied navigation on Goat-Bench [16]: the multi-modal open-vocabulary navigation benchmark. (a) The superiority of our M3DSG over traditional 3D scene graphs. (b) Distance statistics from the goal for the previous method (3D-Mem [41] as an example). (c) Our MSGNav system achieves state-of-the-art performance on Goat-Bench.
@@ 2.1 Graph-based Scene Exploration
Graph representations [11, 17, 35] provide explicit spatial–semantic structures for scene reasoning and integrate well with LLMs/VLMs. Examples include SayPlan [34] for 3D scene graph planning, OVSG [3] for open-vocabulary grounding, imaginative world modeling [13], and BEV graph policies [26]. Graph reasoning has also been applied in navigation, e.g. SG-Nav [42] for hierarchical prompting and graph-retained adaptation [12] for cross-domain VLN. Yet, the overly abstract textual edges of current scene graphs limit the comprehensive scene context understanding, motivating our multi-modal 3D scene graph (M3DSG).
@@ 2.2 Zero-Shot Navigation
Previous embodied navigation methods, primarily based on reinforcement learning (RL) [28, 33], suffer from limited generalization and a significant sim-to-real gap, which hinders their application in real-world environments. Zero-shot navigation allows agents to act in unseen environments without task-specific training, unlike supervised methods [32, 39] which require large-scale simulation. Existing zero-shot navigation methods span object-goal (ON), image-goal (IIN), and text-goal (TN). ON methods include CoW [9] with CLIP [29] frontier exploration, ESC [54], OpenFMNav [20], VLFM [44] using LLM reasoning, UniGoal [43] for multi-modal unification, and RATE-Nav [22] with region-aware termination. IIN methods range from Mod-IIN [19] to SIGN [40] for safety-aware drone navigation. TN methods such as InstructNav [24] generalize instruction following across ON, DDN, and VLN [5, 21, 25, 48, 53]. However, most zero-shot work ignores the last-mile challenge of choosing an optimal viewpoint after target localization.
@@ 3.2.1 M3DSG Structure Definition
During the environment exploring process, M3DSG constructs a scene graph $\mathbf{S}=(\mathbf{O},\mathbf{E})$, where $\mathbf{O}=\{o_{i}\}_{i=1}^{N_{o}}$ and $\mathbf{E}=\{\mathbf{e}_{j}\}_{j=1}^{N_{e}}$ denote the sets of objects and edges, respectively. Here, $N_{o}$ is the number of detected objects, and $N_{e}=\sum_{1\leq x<y\leq N_{o}}\mathds{1}_{\|Pos_{o_{x}}-Pos_{o_{y}}\|<\theta}$ is the number of object pairs $(o_{x},o_{y})$ whose spatial distance is below $\theta$. Each object $o_{i}$ is described by attributes including corresponding category, 3D coordinates, and room locations. Each edge $\mathbf{e}_{j}$ stores a set of RGB-D images $\mathbf{I}_{j}=\{\mathcal{I}_{k}\}_{k=1}^{|\mathbf{I}_{j}|}$ that represent the pair of corresponding objects to record their visual relationship.
@@ 3.2.2 Incremental Construction of M3DSG
Edge Update. The core of an edge is a set of images $\mathbf{I}$ in which an object pair co-occurs, indexed by the unique IDs of the object pair, i.e. $\mathbf{E}:(\mathcal{ID},\mathcal{ID})\rightarrow\mathbf{I}.$
Meanwhile, we maintain a mapping from each image to its corresponding object pair, i.e. $\mathbf{H}:\mathcal{I}\rightarrow(\mathcal{ID},\mathcal{ID}).$
For the current RGB-D frame $\mathcal{I}_{t}$ at step $t$ with object set $\mathbf{O}^{frame}_{t}$ (from Object Update), we generate co-occurring object pairs $(o_{x},o_{y})$ that appear together in $\mathcal{I}_{t}$ as:
@@ 3.3.4 Visibility-based Viewpoint Decision (VVD)
For a navigation agent, the high-level decision process involves the location of the target, while the low-level planner converts it into executable actions. Therefore, an optimal final decision should choose a viewpoint with clear visibility to facilitate subsequent operations. Previous zero-shot methods [10, 15, 41, 42] often assume that knowing the target location is enough, and select the nearest traversable position as the navigation viewpoint. However, as shown in Fig. 4(a), this may yield viewpoints with poor visibility, causing failure even when the target $\bar{o}$ is correctly localized.
We term this the “last-mile” problem of navigation—identifying a suitable viewpoint despite correctly locating the target $\bar{o}$. To solve this problem, we realize that a good viewpoint should have clear visibility of the target. To achieve this goal, we propose a Visibility-based Viewpoint Decision (VVD) module (in Algorithm 2). As shown in Fig. 4(b), VVD uniformly samples candidate viewpoints around target $\bar{o}$ within radii $\mathbf{R}=\{r_{j}\}_{j=1}^{N_{R}}$. For each candidate $\mathbf{v}_{i}$, VVD computes a visibility score by evaluating occlusion between $\mathbf{v}_{i}$ and target point cloud $\mathcal{PC}_{\bar{o}}$. The viewpoint with the highest visibility score $\textbf{v}_{best}$ is selected as the navigation coordinate. Compared with traditional choices, VVD yields viewpoints with superior target visibility. Further explanation and analysis of the VVD module will be supplemented in the Appendix.
@@ 4.1 Experimental Setting
We evaluate our proposed approach on two established goal-oriented navigation benchmarks: GOAT-Bench [16], a multi-modal lifelong open-vocabulary dataset, and the widely adopted HM3D-ObjNav [31] (Habitat 2022 ObjectNav Challenge). To validate the generalization capability of our method to novel environments, all evaluations on GOAT-Bench are conducted on the Val Unseen split. Following standard practice, we assess navigation performance using Success Rate ($\text{SR}=\frac{N_{\text{success}}}{N_{\text{total}}}$) and Success weighted by Path Length ($\mathrm{SPL}=\frac{1}{N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}S_{i}\frac{l^{s}_{i}}{\max(l^{s}_{i},l^{a}_{i})}$), where $S_{i}\in\{0,1\}$ denotes binary task success, and $l^{s}_{i}$ and $l^{a}_{i}$ represent the shortest path distance and the actual path traversed by the agent.
@@
We employ GPT-4o (version 2024-08-06) as our primary Vision-Language Model (VLM) via the OpenAI API. Additional analyses utilizing Qwen-VL-Max as an alternative VLM backbone, along with extended evaluations on the HM3D-OVON benchmark [45], are deferred to the supplementary material.
@@ 4.2.2 HM3D-ObjNav Benchmark
Table 1 demonstrates the superior zero-shot navigation capabilities of our MSGNav on multi-modal open-vocabulary tasks in GOAT-Bench, we further evaluate it on the widely used HM3D-ObjNav benchmark.
As shown in Table 1, MSGNav achieves a state-of-the-art Success Rate (SR) of 63.0%, outperforming the previous best training-free method, Apexnav [51], by 3.4%. While our Success Path Length (SPL) of 31.4% is slightly lower than that of Apexnav, this may be caused by the VVD module prioritizing high-visibility viewpoints over absolute shortest paths, which slightly increases the path length but significantly boosts the overall success rate.
@@
ProcTHOR-ZS [8] × 20.2 8.2
ZSON [27] × 25.5 12.6
PSL [37] × 42.4 19.2
L3MVN [46] 48.7 23.3
GOAT [4] × 50.6 24.1
ESC [54] 39.2 22.3
VLFM [44] 52.4 30.3
OpenFMNav [20] 52.5 24.1
SG-Nav [42] 54.0 24.9
Uni-goal [43] 54.5 25.1
Trihelper [50] 56.5 25.3
Apexnav [51] 59.6 33.0
MSGNav (Ours) 63.0 31.4
@@ 5 Conclusion
In this paper, we propose the MSGNav, a zero‑shot embodied navigation framework built upon a Multi-modal 3D Scene Graph (M3DSG) that preserves visual information for efficient construction, robust perception, and unconstrained vocabulary. Meanwhile, we integrate a visibility‑based viewpoint decision module in MSGNav to resolve the last‑mile problem. Extensive experiments on GOAT‑Bench and HM3D‑OVON datasets demonstrate consistent and notable gains over existing methods, highlighting the effectiveness of the proposed multi‑modal scene representations in embodied navigation.
Limitations and future work. Despite these advantages of MSGNav, scene graph-based methods still face low inference efficiency due to the latency of VFMs and VLMs, suggesting future work on faster graph construction and inference for real‑time deployment.
Acknowledgements. This work was supported in part by the National Natural Science Foundation of China (No.42571514).
@@ 6.3 VFMs for Constructing M3DSG
VLM Overall (2669) Object Category (991) Language (856) Image (822)
GPT-4o [14] 51.97 29.56 53.38 29.98 47.43 25.08 54.99 33.71
Qwen-VL-Max [1] 52.79 30.79 56.10 31.53 45.56 27.18 56.33 33.67
7 More Experimental Results
7.1 Additional Analysis on HM3D-OVON
@@ 8.3 Visualization about exploration
8.4 Analysis of multi-modal scene Graph reasoning

### v4

@@ 1 Introduction
Embodied navigation is a fundamental capability of robotic agents, enabling them to accomplish diverse tasks in realistic environments [47]. Such tasks are typically goal-oriented, with targets defined by objects, natural language descriptions, or images [45]. Recently, the challenging open-vocabulary embodied navigation task [52] has gained increasing attention, as real-world agents are required to generalize to a wide and unconstrained range of object categories [47].
Previous RL-based embodied navigation methods suffer from poor generalization and a large sim-to-real gap [44]. Inspired by rapid advances in pre-trained large model techniques [17, 1], zero-shot embodied navigation approaches have emerged at this historic moment. Zero-shot navigation system does not require any training or fine-tuning before being applied to real-world scenarios. These approaches [43, 18, 53] substantially reduce training time and computational cost while enabling open‑vocabulary navigation, thereby facilitating deployment in real‑world applications.
Notably, a series of methods [36, 37] that build explicit 3D scene graphs from observations and leverage LLMs to drive exploration have shown promising performance on standard goal-oriented navigation benchmarks [41, 2]. However, traditional 3D scene graphs overly abstract 3D object relations into simple textual labels (e.g., “top”, “beside”), which seriously impedes progress on zero-shot embodied navigation. As illustrated in Fig. 1, the extreme abstracted 3D object-relation in traditional 3D scene graphs introduces three main limitations:
@@
Figure 2: Performance comparisons between our MSGNav and other existing methods for embodied navigation on Goat-Bench [19]: the multi-modal open-vocabulary navigation benchmark. (a) The superiority of our M3DSG over traditional 3D scene graphs. (b) Distance statistics from the goal for the previous method (3D-Mem [43] as an example). (c) Our MSGNav system achieves state-of-the-art performance on the challenging Goat-Bench.
@@ 2.1 Graph-based Scene Exploration
Unlike traditional feature representations [16, 14, 15], graph representations [9, 20] provide explicit spatial–semantic structures for scene reasoning and integrate well with LLMs/VLMs. Examples include SayPlan [35] for 3D scene graph planning, OVSG [3] for open-vocabulary grounding, imaginative world modeling [13], and BEV graph policies [30]. Graph reasoning has also been applied in navigation, e.g. SG-Nav [44] for hierarchical prompting and graph-retained adaptation [12] for cross-domain VLN. Yet, the overly abstract textual edges of current scene graphs limit the comprehensive scene context understanding, motivating our multi-modal 3D scene graph (M3DSG).
@@ 2.2 Zero-Shot Navigation
Zero-shot navigation allows agents to act in unseen environments without task-specific training, unlike supervised methods [33, 38] which require large-scale simulation. Existing zero-shot navigation methods span object-goal (ON), image-goal (IIN), and text-goal (TN). ON methods include CoW [7] with CLIP [32] frontier exploration, ESC [51], OpenFMNav [23], VLFM [46] using LLM reasoning, UniGoal [45] for multi-modal unification, and RATE-Nav [25] with region-aware termination. IIN methods range from Mod-IIN [22] to SIGN [42] for safety-aware drone navigation. TN methods such as InstructNav [28] generalize instruction following across ON, DDN, and VLN [4, 24, 29, 49]. However, most zero-shot work ignores the last-mile challenge of choosing an optimal viewpoint after target localization.
@@ 3.2.1 M3DSG Structure Definition
During the environment exploring process, M3DSG constructs a scene graph $\mathbf{S}=(\mathbf{O},\mathbf{E})$, where $\mathbf{O}=\{o_{i}\}_{i=1}^{N_{o}}$ and $\mathbf{E}=\{\mathbf{e}_{j}\}_{j=1}^{N_{e}}$ denote the sets of objects and edges, respectively. Here, $N_{o}$ is the number of detected objects, and $N_{e}=\sum_{1\leq x<y\leq N_{o}}\mathds{1}_{\|Pos_{o_{x}}-Pos_{o_{y}}\|<\theta}$ is the number of object pairs $(o_{x},o_{y})$ whose spatial distance is below $\theta$. Each object $o_{i}$ is described by attributes including corresponding category, 3D coordinates, and room locations. Each edge $\mathbf{e}_{j}$ stores a set of RGB-D images $\mathbf{I}_{j}=\{\mathcal{I}_{k}\}_{k=1}^{|\mathbf{I}_{j}|}$ that represent the pair of corresponding objects to record the visual cues.
@@ 3.2.2 Incremental Construction of M3DSG
Edge Update. The core of an edge is a set of images $\mathbf{I}$ in which an object pair co-occurs, indexed by the unique IDs of the object pair, i.e. $\mathbf{E}:(\mathcal{ID},\mathcal{ID})\rightarrow\mathbf{I}.$ Meanwhile, we maintain a mapping from each image to its corresponding object pair, i.e. $\mathbf{H}:\mathcal{I}\rightarrow(\mathcal{ID},\mathcal{ID}).$ For the current RGB-D frame $\mathcal{I}_{t}$ at step $t$ with object set $\mathbf{O}^{frame}_{t}$ (from Object Update), we generate co-occurring object pairs $(o_{x},o_{y})$ that appear together in $\mathcal{I}_{t}$ as:
@@ 3.3.4 Visibility-based Viewpoint Decision (VVD)
For a navigation agent, the high-level decision process involves the location of the target, while the low-level planner converts it into executable actions. Therefore, an optimal final decision should choose a viewpoint with clear visibility to facilitate subsequent operations [40, 39]. Previous zero-shot methods [8, 18, 43, 44] often assume that knowing the target location is enough, and select the nearest traversable position as the navigation viewpoint. However, as shown in Fig. 4(a), this may yield viewpoints with poor visibility, causing failure even when the target $\bar{o}$ is correctly localized.
We term this the “last-mile” problem of navigation—identifying a suitable viewpoint despite correctly locating the target $\bar{o}$. To solve this problem, we realize that a good viewpoint should have clear visibility of the target. To achieve this goal, we propose a Visibility-based Viewpoint Decision (VVD) module (in Algorithm 2). As shown in Fig. 4(b), VVD uniformly samples candidate viewpoints around target $\bar{o}$ within radii $\mathbf{R}=\{r_{j}\}_{j=1}^{N_{R}}$. For each candidate $\mathbf{v}_{i}$, VVD computes a visibility score by evaluating occlusion between $\mathbf{v}_{i}$ and target point cloud $\mathcal{PC}_{\bar{o}}$. The viewpoint with the highest visibility score $\textbf{v}_{best}$ is selected as the navigation coordinate. Compared with traditional choices, VVD yields viewpoints with superior target visibility.
@@ 4.1 Experimental Setting
We evaluate our proposed approach on two established goal-oriented navigation benchmarks: 1) GOAT-Bench [19] (Multi-modal lifelong open-vocabulary dataset, 360 episodes, 36 scenes, 2669 total subtasks, 36 novel goal categories). 2) HM3D-ObjNav (HM3D-Semantics-v0.2 [41] from 2023 Habitat Challenge, 1000 episodes, 36 scenes, 6 goal categories). To validate the generalization capability of our method to novel environments, all evaluations on GOAT-Bench are conducted on the Val Unseen split. Following standard practice, we assess navigation performance using Success Rate ($\text{SR}=\frac{N_{\text{success}}}{N_{\text{total}}}$) and Success weighted by Path Length ($\mathrm{SPL}=\frac{1}{N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}S_{i}\frac{l^{s}_{i}}{\max(l^{s}_{i},l^{a}_{i})}$), where $S_{i}\in\{0,1\}$ denotes binary task success, and $l^{s}_{i}$ and $l^{a}_{i}$ represent the shortest path distance and the actual path traversed by the agent.
@@
We used the official success distance threshold from the benchmark: 0.25 for GoatBench and 1.0 for HM3D-ObjNav. We employ GPT-4o (2024-08-06) as our primary Vision-Language Model (VLM) via the OpenAI API. Additional analyses utilizing Qwen-VL-Max as an alternative VLM backbone, along with extended evaluations on the HM3D-OVON benchmark [47], are deferred to the supplementary material.
@@ 4.2.2 HM3D-ObjNav Benchmark
Table 1 demonstrates the superior zero-shot multi-modal open-vocabulary navigation capabilities of our MSGNav, we further evaluate it on the widely used HM3D-ObjNav benchmark. As shown in Table 1, MSGNav achieves a state-of-the-art Success Rate (SR) of 74.1%, which is 1.9% higher than that of the previous best-performing method WMNav [31], and significantly outperforms other prior methods. Although our Success Path Length (SPL) is nearly the same as WMNav without any significant advantage, this may be because the VVD module prioritizes viewpoints with a wide field of view over the absolute shortest path.
@@
L3MVN [48] 36.3 15.7
SG-Nav [44] 49.6 25.5
InstructNav [28] 58.0 20.9
CompassNav [26] 59.6 26.9
Schrodinger’sNav [11] 60.9 23.7
VLFM [46] × 62.6 31.0
DORAEMON [10] 66.5 20.6
WMNav [31] 72.2 33.3
MSGNav (Ours) 74.1 33.4
@@ 5 Conclusion
In this paper, we propose the MSGNav, a zero‑shot embodied navigation framework built upon a Multi-modal 3D Scene Graph (M3DSG) that preserves visual information for efficient construction, robust perception, and unconstrained vocabulary. Meanwhile, we integrate a visibility‑based viewpoint decision module in MSGNav to resolve the last‑mile problem. Extensive experiments on GOAT‑Bench and HM3D‑ObjNav datasets demonstrate consistent and notable gains over existing methods, highlighting the effectiveness of the proposed multi‑modal scene representations in embodied navigation.
Limitations and future work. Despite these advantages of MSGNav, scene graph-based methods still face low inference efficiency due to the latency of VFMs and VLMs, suggesting future work on faster graph construction and inference for real‑time deployment. Additionally, while the “last-mile” problem has been mitigated by the VVD module, it is not fully resolved. Exploring reinforcement learning approaches, e.g. active perception, to further solve this challenge is also a way worthy of attention.
Acknowledgements. This work was supported in part by the National Natural Science Foundation of China (No. 42571514) and the Zhongguancun Academy (Grant No. 02012405).
@@ 6.3 VFMs for Constructing M3DSG
Figure 7: Detailed end-to-end processing flow of the proposed framework. The system executes a continuous sense-reason-act loop across four key stages: (1) multi-modal 3D scene graph construction via Vision Foundation Models (VFMs), (2) closed-loop reasoning based on the extracted sub-graph $\mathbf{S}^{k}$ and historical memory, (3) adaptive updates to the semantic vocabulary and decision memory, and (4) visibility-based viewpoint decision or frontier exploration. Please refer to the corresponding text for a detailed description of each mathematical symbol and data flow.
@@
6.4 Illustration of End-to-end Framework
In addition to the framework diagram in the main paper, we also provide a more intuitive end-to-end framework diagram that systematically illustrates the data flow and processing logic between modules, as shown in Fig. 7.
The system operates in a continuous sense-reason-act loop at each timestep $t$, which can be detailed in four key stages: (1) Constructing Multi-modal 3D Scene Graph: Given the current observation $\mathcal{I}_{t}$ and agent pose, Vision Foundation Models (VFMs) extract visual, spatial, and room-level attributes to incrementally update the object nodes and relational edges of the scene graph $\mathbf{S}_{t}$. (2) Effective and Closed-Loop Reasoning: A Key Sub-graph Selection (KSS) module filters $\mathbf{S}_{t}$ into a task-relevant sub-graph $\mathbf{S}^{k}$. An LLM agent then reasons over $\mathbf{S}^{k}$, historical decision memory $M_{t-1}$, and goal information $\mathcal{G}$ to generate an action response $R_{t}$ and proposed vocabulary updates $\tilde{V}_{t}$. (3) Vocabulary and Decision Update: The Adaptive Vocabulary Update (AVU) module integrates $\tilde{V}_{t}$ to form the current vocabulary $V_{t}$, while the Closed-Loop Reasoning (CLR) module records $R_{t}$ to update the decision memory $M_{t}$. (4) Agent Step with Visibility-based Viewpoint Decision: If the target is identified, the VVD module selects the optimal viewpoint $\mathbf{v}_{best}$ that maximizes the visibility score of the target point cloud $PC_{\bar{o}}$; otherwise, the LLM selects a new frontier for exploration in step $t+1$.
VLM Overall (2669) Object Category (991) Language (856) Image (822)
GPT-4o [17] 51.97 29.56 53.38 29.98 47.43 25.08 54.99 33.71
Qwen-VL-Max [1] 52.79 30.79 56.10 31.53 45.56 27.18 56.33 33.67
7 More Experimental Results
7.1 Additional Analysis on HM3D-OVON
@@ 8.3 Visualization about Exploration
8.4 Analysis of Reasoning based on M3DSG